OpenStudy (anonymous):

WILL FAN AND MEDAL What are the explicit equation and domain for a geometric sequence with a first term of 4 and a second term of -8?

OpenStudy (anonymous):

an=4(-2)^n-1; all integers where n≥0 an=4(-2)^n-1; all integers where n≥1 an=4(-12)^n-1; all integer where n≥0 an=4(-12)^n-1; all integer where n≥1

OpenStudy (campbell_st):

ok... just a correct.... you answer choices are for a geometric sequence and your question mentions an arithmetic sequence..

OpenStudy (anonymous):

Sorry fixed it

OpenStudy (campbell_st):

so to get from 4 to -8 by multiplying the common ratio is -2 and a term in a geometric sequence is found using \[a_{n} = a_{1} \times r^{n - 1}\] where \[a_{1} ~ 1st ~term~~~~~~and~~~ r = common ~ratio\] the domain is \[n \ge 1\] so substitute what you know into the general equation to get the explicit form

OpenStudy (anonymous):

Oh okay thank you :)